# Brain states analysis of EEG predicts multiple sclerosis and mirrors disease duration and burden

### v2

@@ Abstract
Objectives: The proposed method deciphers electrical brain states which as multi-dimensional cognetoms quantitatively discriminate normal from pathological patterns in an EEG.
Methods: Baseline recordings from a prior EEG study of 93 subjects, 37 with MS, were analyzed. Spectral bands served to compute cognetoms and categorize subsequent feature combination sets.
Results: A significant correlation arose between brain states predictors, clinical data and disease duration. Using cognetoms and spectral bands, a cross-sectional comparison separated patients from controls with a precision of 82% while using bands alone arrived at 64%.
Conclusions: Brain states analysis successfully distinguishes controls from patients with MS. The congruity with disease duration is a neurobiological indicator for disease accumulation over time. Our results imply that data-driven comparisons of EEG data may complement customary diagnostic methods in neurology and psychiatry. However, thinking ahead for quantitative monitoring of disease time course and treatment efficacy, we hope to have established the analytic principles applicable to longitudinal clinical studies.
@@ 1 Introduction
Mental illness is an enormous burden for societies worldwide.Smith (2011); Vigo et al. (2016) Above all it is the decay of cognitive functioning, the essence of human reasoning, that turns a mental illness into a devastating, painful and tangible experience for patient and surrounding alike. Nonetheless diagnosis and staging of diseases still rely on customary clinical, laboratory and imaging tests by which cognitive integrity is not assessed on biological grounds. Electrical scalp recordings exist for a century and were used decades ago as diagnostic evidence in multiple sclerosis although the analytic methods lacked the quantitative aspect for relating brain measures to disease time course or treatment outcome. We believe that a modern data-driven functional brain mapping analysis approach using non-invasive EEG data will meaningfully enhance the prognostic power for clinical trials and facilitate the routine assessment of drug efficacy and safety through objective measures for a changing cognito-electrical substrate.
@@
The central nervous system (CNS) is an electrical organ and invokes with each mental act a plethora of neural activities instantly reflected in the electrical brain space and recordable on the skull surface with electrode arrays.Buzsáki (2005) Electrical states of brain functioning, coined cognetoms, change at any given momentKoenig et al. (1999); Samdin et al. (2017) and are traced with physiological measures while anatomical brain lesions evolve slower over days or weeks and can be monitored with MRI exams.Eshaghi et al. (2021); Tommasin et al. (2021); Lazzarotto et al. (2024) The crucial difference is in the immediacy and temporal accuracy of the electrical measures and the dispersed soft demarcation of the electrical phenomena on the skull surface that mirror underlying distributed brain activity. These diffuse and perpetually evolving cloud-like arrangements are due to oscillating neural networks and reflect equally housekeeping and cognitive neural processes.Buzsáki (2005)
@@
MS is a chronic, relapsing-remitting, inflammatory disease with multiple foci of demyelination disseminated over space and time throughout the CNS. It can lead to impaired cognitive reserves and capacity,Maggi et al. (2025) disabilityCobo-Calvo et al. (2025) and early death. Northern countries show a high frequency for MS with the highest prevalence in Canada of 0.3%.Amankwah et al. (2017)
@@
Recent advances in pharmaceutical research for aggressive disease-modifying treatments lead to a surge in the development of novel drugs for MS over the last two decades. Currently, there are 18 disease-modifying therapies approved by Health Canada (mostly for the relapsing remitting form) while heaps of new compounds are being tested in clinical trials for clinical and MRI outcomes. The desired effect of drug therapy on cognitive preservationFalet et al. (2022) however is still conventionally assessed with questionnaires and neuro-psychological tests although a recent upswing of publications about MS using functional brain measures has been observed.Kisler et al. (2020); Paolicelli et al. (2021); Santinelli et al. (2021); Mohseni and Moghaddasi (2022); Sorrentino et al. (2022)
@@
We adopted the statistical principles of our former brain states analysis methodJanoos et al. (2011); Mórocz et al. (2012); Janoos et al. (2013) and tested the new method on data taken from a prior EEG study about multiple sclerosis in healthy subjects and patients.Jouzizadeh et al. (2021) Our motivation to analyze such cross-sectional cohort data was rooted in the current lack of access to serial EEG measures in MS patients. While our methods are primarily developed for the purpose of comparing serial EEG measures longitudinally within the same individual we felt confident that the method’s data modeling phase could also generate robust enough brain state maps from single-time-point data to meaningfully support a cross-sectional comparison of brain states features between two subject cohorts. We show here the successful tangential application of a brain states EEG analysis method that distinguishes healthy data from MS data.
@@ 2.1 Subjects and EEG recording
The EEG data of 93 subjects were taken from an earlier project where participant criteria and recording sessions have been described in detail in the original publication.Jouzizadeh et al. (2021) Participants signed consent forms for all procedures and data analyses in accordance with the Declaration of Helsinki. The anonymized data of 37 subjects diagnosed with MS (24 female and 13 male, mean age 29.5 years, age range 20-53 years) and 56 healthy volunteers (29 female and 27 male, mean age 20.7 years, age range 18-31 years) were taken into consideration for this data analysis project. The mean of the Expanded Disability Status Scale (EDSS) scores for MS patients at EEG acquisition time was 1.96 (EDSS score range 0-7). The diagnosis of MS was based on the revised McDonald criteria for relapsing-remitting MS (RRMS) as used at Boualicina Hospital, Sari, Iran. The mean disease duration after confirmed diagnosis was at the time of the EEG recording 5.1 years (range 0-18 years). Resting-state EEG was recorded for 5 minutes for each condition, eyes-open (EO) and eyes-closed (EC), using a Brainmaster Discovery24 data acquisition system and 19 scalp electrode positions at a sampling rate of 256 Hz.
@@ 2.2 Data preparation
Raw binary sensor data were extracted as time stream vectors from original EEG data files sans further metadata or personal information. Electrode labels and anatomical positional information were discarded and hence played no further role during data analysis. Attention was paid to maximally preserve the fine-grained intrinsic physiological information present in the data Meisler et al. (2019) for the brain states generation phase further downstream in the processing pipeline. For these reasons EEG data were taken as is without applying common data filtering, cleansing and artifact removal techniques for noise reduction, faulty electrodes, eye blink, cardiac or other muscle activity, and hence without introducing pre-processing related new artifacts.McDermott et al. (2021) This avenue is possible because our method of data processing during brain state detection and later analysis phases intrinsically suppresses noise components in the data throughout the entire brain states processing pipelineJanoos et al. (2011); Mórocz et al. (2012) and renders the method fundamentally apt and robust in dealing with clinical-grade real-life data.Janoos et al. (2013) This ‘automatic noise robustness’ lies in the fact that brain states detected throughout the data sets bear inter-individual characteristics and are hence not prone to include meaningless individual noise properties not universally present across the entire data set.
@@ 2.3 Brain states detection (BSD)
We define brain states, or cognetoms, as snapshots of repeatable patterns of highly correlated signals. Brain states are understood as moments of distributed brain activity of a given time point.Koenig et al. (1999); Janoos et al. (2011); Mórocz et al. (2012) In a first move a limited set of eight brain states are computed. These sets serve as lowest-level elements, or level-0 cognetoms, quasi as alphabet letters or building blocks to use in future aggregation cascade models of ever increasing cognetom complexity and abstraction level. EEG recordings were segmented into data epochs where each individual epoch, a brain state, consists of 500 EEG frames (scan frequency at 256Hz, epoch length 1.953sec), and are described by five EEG frequency band amplitudes (delta: 0-4Hz, theta: 4-8Hz, alpha: 8-14Hz, beta : 14-32Hz, gamma: 32-45Hz) that are extracted from raw EEG data.
@@ 3.1 Brain states analysis (BSA)
In order to capture differences between the EEG data of healthy persons and patients with MS two approaches were applied : manifold learning and catboost classifier. For this, 80 distinct feature combinations were built and are described as combinations of 8 brain state weights, two experimental recording conditions (eyes-open and eyes-closed) and the five EEG frequency band amplitudes as mentioned earlier.
The first approach, manifold learning, reduces the dimensionality of the individual data.McInnes et al. (2018) The previously prepared sequence of brain states and weights are used to embed the individual data into a two-dimensional space. Each point in that space corresponds with one person. First we applied the non-supervised version of the UMAP algorithm for which the results reveal a distinct trend separating control from MS cases (upper figure 1). Next we applied the supervised version by splitting data into train (60%) and test (40%) subsets of data. The trained model produced the two-dimensional embedding with high separability between the two subject groups (lower figure 1).
Our second approach used the catboost classifier methodDorogush et al. (2018) which first splits the data into train (60%) and test (40%) subsets of data, then trains the model, and last evaluate the quality of the model on the test data subset. The quality of the model is described by precision and F-score values.
@@ 3.3 Specific feature combinations for MS detection
The relative importance of 80 feature combinations (8 brain states x 5 frequency bands x 2 experimental conditions) that separate the two subject groups into healthy vs MS is outlined in the graphs of figure 2. We observe that a set of certain feature combinations that have a high relative importance, include both eyes-open and eyes-closed conditions and the delta frequency band appear to drive the separation between healthy and MS subjects.
@@ 3.4 Disease predictor compared to clinical data
Both plots in figure 3 imply a significant correlation of these two clinical measures with the brain states predictor space, which, we surmise, separates the two subject groups, the more the disease burden in the CNS increases. In fact, all patients to the far left in the ‘control’ brain states predictor space (marked cyan with letters from a to g) feature a short disease history of less than two years (mean 0.8 years) and cited motor and sensory deficits which may originate from lesions in the spinal cord rather than the upper brain structures. In contrast, the patients to the right (marked in red) with longer disease histories (mean 6.0 years), self-reported also vision, balance and brain stem disturbances beside only motor and sensory issues in the limbs. This suggests more complex neurological deficit patterns that are likely related to an increase of lesion load in the cerebrum itself, the de-facto generator for the EEG signal.
Figure 3: Distinction of control subjects from patients with MS through brain states prediction. The disease predictor value corresponds to the rotated coordinates as seen in Fig. 1B. A rotation was performed using the principal components of the point distribution where the predictor value is considered as the primary principal component. (A) The upper graph illustrates this separability plotted against EDSS measures that were obtained in MS. Patients are shown as red or cyan dots while controls feature blue dots. Seven patients are found as outliers on the left side of the disease predictor axis and are marked with cyan characters (a-g) for further discussion under Results. A linear regression is shown as gray line. (B) The lower graph plots separability against disease duration in years. Two patients (d,f) lack data on disease duration and hence are not included here. Linear regression is shown as gray trend line.
@@ 4 Discussion
We successfully transformed a brain states hypothesisKoenig et al. (1999); Janoos et al. (2011) into an analytic method for processing EEG data using a quantitative embedding and modeling approach for brain states and applied the new technique to separate the two subject groups of an earlier EEG study about MS. We observed several beneficial aspects when brain states analysis is added in a context as described here.
@@ 4.1 Implications
First we demonstrate with an initial implementation of our novel algorithm as a two-phase processing pipeline with brain states detection and brain states analysis that already regular clinical EEG data clearly carry the necessary information to discriminate two subjects groups as shown here. This is to our knowledge a first successful demonstration of such a hypothesis-free space-agnostic data-driven full-brain raw-data EEG analysis approach for MS. Second we state that the inclusion of brain states in the analysis is sensible and significantly enhances the statistical robustness of the approach. Third we believe this brain states approach is easily generalizable to other disease conditions and importantly also for the intra-individual monitoring of serially acquired longitudinal EEG data.
The current method readily distinguished healthy from MS cases using clinical EEG data in spite of obstacles faced with such as i) early stage of the software implementation, ii) swap in the original purpose of the approach from monitoring serial data to comparing subjects, and iii) low information density in common clinical EEG recordings. We explain the analysis outcome with flow derangement between electrical mental processes in MS secondary to altered physical brain properties.Kisler et al. (2020); Paolicelli et al. (2021) This was newly confirmed in a recent study about disturbances of conductance velocity through white matter lesions.Sorrentino et al. (2022) We currently offer no method to quantify processing speed in brain states but we know already that certain feature combinations of high importance carry slow delta EEG components associated with both eye-open and eyes-closed conditions that appear to drive currently the separation of controls from patients (see figure 2).
We speculate that even just small conductance alterations in MS will perceptibly skew the ’electric symphony’ among the many brain activities, ’handshakes’, synchronicity and interplay between resonating loops,Buzsáki (2005) all the while the patient still compensates without visible clinical deficit. However, our concept about internal clockworks being disturbed between brain activities will remain hypothetical without having access to high-quality EEG data.
@@ 4.2 Significance
That this method succeeds on clinical-grade EEG data rests on two facts, in the processing of complete brain states and the characterizing of entire EEG recordings in form of comprehensive brain states sequences without cleansing noisy epochs. These two points highlight potential advantages of the method in that older less-advanced EEG databases could be sensibly re-analyzed with modern brain states analysis techniques, while collecting new EEG data in MS or other conditions could for budgetary reasons happen on affordable low-resolution EEG recording devices. Having said that, only high sensor density, high scan frequency and rich cognitive stimulative task conditions will leverage dense enough information in the EEG and hence contribute to deep brain states abstraction models. While clinical EEG data come in their own right only the highest-quality EEG data can deliver the fodder to study the internal syntax inside multi-hierarchical cognetom cascade models.

### v5

@@ Abstract
Objective: The proposed method deciphers electrical brain states which as multi-dimensional cognetoms quantitatively discriminate normal from pathological patterns in an EEG.
Method: Baseline recordings from a prior EEG study of 93 subjects, 37 with MS, were analyzed. Spectral bands served to compute cognetoms and categorize subsequent feature combination sets.
Result: A significant correlation arose between brain states predictors, clinical data and disease duration. Using cognetoms and spectral bands, a cross-sectional comparison separated patients from controls with a precision of 82% while using bands alone arrived at 64%.
Conclusion: Brain states analysis successfully distinguishes controls from patients with MS. The congruity with disease duration is a neurobiological indicator for disease accumulation over time. Our results imply that data-driven comparisons of EEG data may complement customary diagnostic methods in neurology and psychiatry. However, thinking ahead for quantitative monitoring of disease time course and treatment efficacy, we hope to have established the analytic principles applicable to longitudinal clinical studies.
@@ 1 Introduction
Mental illness is an enormous burden for societies worldwide (Smith, 2011; Vigo et al., 2016). Above all it is the decay of cognitive functioning, the essence of human reasoning, that turns a mental illness into a devastating and tangible experience for patient and surrounding alike. Nonetheless diagnosis and staging of diseases still rely on customary clinical, laboratory and imaging tests by which cognitive integrity is not assessed on biological grounds. Electrical scalp recordings exist for a century and were used decades ago as diagnostic evidence in multiple sclerosis although the analytic methods lacked the quantitative aspect for relating brain measures to disease time course or treatment outcome. We believe that a modern data-driven functional brain mapping analysis approach using non-invasive EEG data will meaningfully enhance the prognostic power for clinical trials and facilitate the routine assessment of drug efficacy and safety through objective measures for a changing cognito-electrical substrate.
@@
The central nervous system (CNS) is an electrical organ and invokes with each mental act a plethora of neural activities instantly reflected in the electrical brain space and recordable on the skull surface with electrode arrays (Buzsáki, 2005). Electrical states of brain functioning, coined cognetoms, change at any given moment (Koenig et al., 1999; Samdin et al., 2017) and are traced with physiological measures while anatomical brain lesions evolve slower over days or weeks and are monitored with MRI exams (Eshaghi et al., 2021; Tommasin et al., 2021; Lazzarotto et al., 2024). The crucial difference is in the immediacy and temporal accuracy of electrical measures and the dispersed soft demarcation of electrical phenomena on the skull surface that mirror underlying distributed brain activity. These diffuse and perpetually evolving cloud-like arrangements are due to oscillating neural networks and reflect equally housekeeping and cognitive neural processes (Buzsáki, 2005).
@@
MS is a chronic, relapsing-remitting, inflammatory disease with multiple foci of demyelination disseminated over space and time throughout the CNS. It can lead to impaired cognitive reserves and capacity (Maggi et al., 2025) disability (Cobo-Calvo et al., 2025) and early death. Northern countries show a high frequency for MS with the highest prevalence in Canada of 0.3% (Amankwah et al., 2017).
@@
Recent advances in pharmaceutical research for aggressive disease-modifying treatments lead to a surge of novel drugs for MS over the last two decades. Currently, there are 18 disease-modifying therapies approved by Health Canada (relapsing remitting form) while heaps of new compounds are being tested in clinical trials for clinical and MRI outcomes. The desired effect of drug therapy on cognitive preservation (Falet et al., 2022) however is still conventionally assessed with questionnaires and neuro-psychological tests although a recent upswing of publications using functional brain measures in MS is observable (Kisler et al., 2020; Paolicelli et al., 2021; Santinelli et al., 2021; Mohseni and Moghaddasi, 2022; Sorrentino et al., 2022).
@@
We adopted the statistical principles of our former brain states analysis method (Janoos et al., 2011; Mórocz et al., 2012; Janoos et al., 2013) and tested the new method on data taken from a prior EEG study about multiple sclerosis in healthy subjects and patients (Jouzizadeh et al., 2021). Our motivation to analyze such cross-sectional cohort data roots in the current lack of access to serial EEG measures in MS patients. While our methods are primarily developed for the purpose of comparing serial EEG measures longitudinally within the same individual we felt confident that the data modeling phase could also generate robust enough brain state maps from single-time-point data to meaningfully support a cross-sectional comparison of brain states features between two subject cohorts. We show here the tangential application of a brain states EEG analysis method that successfully distinguishes healthy data from MS data.
@@ 2.1 Subjects and EEG recording
The EEG data of 93 subjects were taken from an earlier project where participant criteria and recording sessions have been described in detail in the original publication (Jouzizadeh et al., 2021). Participants signed consent forms for all procedures and data analyses in accordance with the Declaration of Helsinki. The anonymized data of 37 subjects diagnosed with MS (24 female, 13 male, mean age 29.5 years, age range 20-53 years) and 56 healthy volunteers (29 female, 27 male, mean age 20.7 years, age range 18-31 years) were taken into consideration for this data analysis project. The mean of the Expanded Disability Status Scale (EDSS) scores for MS patients at EEG acquisition time was 1.96 (EDSS score range 0-7). The diagnosis of MS was based on the revised McDonald criteria for relapsing-remitting MS (RRMS) as used at Boualicina Hospital, Sari, Iran. The mean disease duration after confirmed diagnosis was at EEG recording time 5.1 years (range 0-18 years). Resting-state EEG was recorded for 5 minutes for each condition, eyes-open (EO) and eyes-closed (EC), using a Brainmaster Discovery24 data acquisition system and 19 scalp electrode positions at 256 Hz sampling rate.
@@ 2.2 Data preparation
Raw binary sensor data were extracted as time stream vectors from original EEG data files sans further metadata or personal information. Electrode labels and anatomical positional information were discarded and played no further role during data analysis. Attention was paid to maximally preserve the fine-grained intrinsic physiological information present in the data (Meisler et al., 2019) for the brain states generation phase further downstream in the processing pipeline. For these reasons EEG data were taken as is without applying common data filtering, cleansing and artifact removal techniques for noise reduction, faulty electrodes, eye blink, cardiac or muscle activity, and without introducing pre-processing related artifacts (McDermott et al., 2021). This avenue is possible because the data processing methods during brain state detection and later analysis phases intrinsically suppress noise components in the data throughout the entire brain states processing pipeline (Janoos et al., 2011; Mórocz et al., 2012) and render the approach fundamentally apt and robust in dealing with clinical-grade real-life data (Janoos et al., 2013). This ‘automatic noise robustness’ lies in the fact that brain states detected throughout the data sets bear inter-individual characteristics and are hence not prone to include meaningless individual noise properties not universally present across the entire data set.
@@ 2.3 Brain states detection (BSD)
We define brain states, or cognetoms, as snapshots of repeatable patterns of highly correlated signals. Brain states are understood as moments of distributed brain activity of a given time point (Koenig et al., 1999; Janoos et al., 2011; Mórocz et al., 2012). In a first move a limited set of eight brain states are computed. These sets serve as lowest-level elements, or level-0 cognetoms, quasi as alphabet letters or building blocks to use in future aggregation cascade models of ever increasing cognetom complexity and abstraction level. EEG recordings were segmented into data epochs where each individual epoch, a brain state, consists of 500 EEG frames (scan frequency at 256Hz, epoch length 1.953sec), and are described by five EEG frequency band amplitudes (delta: 0-4Hz, theta: 4-8Hz, alpha: 8-14Hz, beta : 14-32Hz, gamma: 32-45Hz) that are extracted from raw EEG data.
@@ 3.1 Brain states analysis (BSA)
In order to capture differences between the EEG data of healthy persons and patients with MS two approaches were applied : manifold learning and catboost classifier. For this, 80 distinct feature combinations were built and are described as combinations of 8 brain state weights, two experimental recording conditions (eyes-open and eyes-closed) and five EEG frequency band amplitudes as mentioned.
The first approach, manifold learning, reduces the dimensionality of the individual data (McInnes et al., 2018). The previously prepared sequence of brain states and weights are used to embed the individual data into a two-dimensional space. Each point in that space corresponds with one person. First we applied the non-supervised version of the UMAP algorithm for which the results reveal a distinct trend separating control from MS cases (upper figure 1). Next we applied the supervised version by splitting data into train (60%) and test (40%) subsets of data. The trained model produced the two-dimensional embedding with high separability between the two subject groups (lower figure 1).
Our second approach used the catboost classifier method (Dorogush et al., 2018) which first splits the data into train (60%) and test (40%) subsets of data, then trains the model, and last evaluate the quality of the model on the test data subset. The quality of the model is described by precision and F-score values.
@@ 3.3 Specific feature combinations for MS detection
The relative importance of 80 feature combinations (8 brain states x 5 frequency bands x 2 experimental conditions) that separate the two subject groups into healthy vs MS is outlined in figure 2. We observe that a set of certain feature combinations that have a high relative importance, include both eyes-open and eyes-closed conditions and the delta frequency band appear to drive the separation between healthy and MS subjects.
@@ 3.4 Disease predictor compared to clinical data
Both plots in figure 3 imply a significant correlation of these two clinical measures with the brain states predictor space, which, we surmise, separates the two subject groups, the more the disease burden in the CNS increases. In fact, all patients to the far left in the ‘control’ brain states predictor space (marked green with letters from a to g) feature a short disease history of less than two years (mean 0.8 years) and cited motor and sensory deficits which may originate from lesions in the spinal cord rather than the upper brain parts. In contrast, the patients to the right (marked red) with longer disease histories (mean 6.0 years), self-reported also vision, balance and brain stem disturbances beside only motor and sensory issues in the limbs. This suggests more complex neurological deficit patterns that are likely related to an increase of lesion load in the cerebrum itself, the de-facto generator for the EEG signal.
Figure 3: Distinction of control subjects from patients with MS through brain states prediction. The disease predictor value corresponds to the rotated coordinates as seen in Fig. 1B. A rotation was performed using the principal components of the point distribution where the predictor value is considered as the primary principal component. (A) The upper graph illustrates this separability plotted against EDSS measures that were obtained in MS. Patients are shown as red or green dots while controls feature blue dots. Seven patients are found as outliers on the left side of the disease predictor axis and are marked with green characters (a-g) for further discussion under Results. A linear regression is shown as gray line. (B) The lower graph plots separability against disease duration in years. Two patients (d,f) lack data on disease duration and hence are not included here. Linear regression is shown as gray trend line.
@@ 4 Discussion
We successfully transformed a brain states hypothesis (Koenig et al., 1999; Janoos et al., 2011) into an analytic method for processing EEG data using a quantitative embedding and modeling approach for brain states and applied the new technique to separate the two subject groups of an earlier EEG study about MS (Jouzizadeh et al., 2021). We observed several beneficial aspects when brain states analysis is added in a context as described here.
@@ 4.1 Implications
First we demonstrate with an initial implementation of our novel algorithm, a two-phase processing pipeline with brain states detection and brain states analysis, that already regular clinical EEG data clearly carry the necessary information to discriminate two cohorts as shown here. This is to our knowledge a first successful demonstration of such a hypothesis-free space-agnostic data-driven full-brain raw-data EEG analysis approach for MS. Second we state that the inclusion of brain states in the analysis is sensible and significantly enhances the statistical robustness of the approach. Third we believe this brain states approach is easily generalizable to other disease conditions and importantly also for the intra-individual monitoring of serially acquired longitudinal EEG data.
The current method readily distinguished healthy from MS cases using clinical EEG data in spite of obstacles faced with : i) early stage of the software implementation; ii) swap in the original purpose of the approach from monitoring serial data to comparing subjects; iii) low information density in common clinical EEG recordings. We explain the analysis outcome with flow derangement between electrical mental processes in MS secondary to altered physical brain properties (Kisler et al., 2020; Paolicelli et al., 2021). This was newly confirmed in a recent study about disturbances of conductance velocity through white matter lesions (Sorrentino et al., 2022). We currently offer no method to quantify processing speed in brain states but we know already that certain feature combinations of high relevance carry slow delta EEG components associated with both eye-open and eyes-closed conditions that appear to drive the separation of controls from patients (see figure 2).
We speculate that even just small conductance alterations in MS will perceptibly skew the ’electric symphony’ among the many brain activities, ’handshakes’, synchronicity and interplay between resonating loops (Buzsáki, 2005), all the while the patient still compensates without visible clinical deficit. However, our concept about internal clockworks being disturbed between brain activities will remain hypothetical without having access to high-quality EEG data (Formica et al., 2025).
@@ 4.2 Significance
That this method succeeds on clinical-grade EEG data rests on two facts, in the processing of complete brain states and the characterizing of entire EEG recordings in form of comprehensive brain states sequences without cleansing noisy epochs. These two points highlight potential advantages of the method in that older less-advanced EEG databases could be sensibly re-analyzed with modern brain states analysis techniques, while collecting new EEG data in MS or other conditions could for budgetary reasons happen on affordable low-resolution EEG recording devices. Having said that, only high sensor density (Formica et al., 2025), high scan frequency and rich cognitive stimulative task conditions will leverage dense enough information in the EEG and hence contribute to deep brain states abstraction models. While clinical EEG data come in their own right only the highest-quality EEG data can deliver the fodder to study the internal syntax inside multi-hierarchical cognetom cascade models.